A sample of Freon-12 (CF2Cl2) occupies 10.0 L at 312 K and 109.00 kPa. Find its volume at STP.
Answer #1

109.00KPa    = 109*0.00986923atm

                      = 1.07575atm

Experimental conditions                                                              STP conditions

P1 = 109KPa                                                                     P2 = 101.325KPa

V1 = 10L                                                                                    V2 =

T1 = 312K                                                                                   T2 = 273K

                P1V1/T1     =   P2V2/T2

               V2              = P1V1T2/T1P2

                                 = 109*10*273/(312*101.325)

                                 = 9.4L

The volume in STP   = 9.4L >>>answer